OpenStudy (anonymous):

You have $2500 to invest at 6% interest compounded quarterly. For how many years will the money need to be invested for that amount to triple?

OpenStudy (gabylovesyou):

quarterly rate = .06/4 = .015 we want the accumulation factor to equal 3 after n quarters: 1.015^n = 3 n*ln1.015 = ln3 n = ln3/ln1.015 =73.7887 quarters divide by 4 and you get 18.447 years = 18 yrs + .447(12) months = 18 yrs, 5+ months (if you will withdraw the money at the end of a month, wait 18 years, 6 months)